a. 0 b. 1 c. 50 d. 2 33. All of the following might explain a firm offering quantity discounts encerpt a. lower costs of handling large orders. b. an inelastic demand for the good. c. monopoly power in this market. d. existence of some high and some low demand consumers 34. Suppose demand for a good is Qo-100-P and supply is Qs-20+P. What is the amount comsumens py producers? a. 60 b. 2400 С. 3600 d. 6400 Suppose goods Xand Yare produced along a prod ction possibilities for terr-狎-500 and they are perfect substitutes such that U-X+Y. The slope of the production possibilities fromticer 35. 500- What is the MRTS at the optimal point? 0.
Answer #1

a) "B"

If the demand for the good is inelastic then the consumer will not get any discount but they have to pay a higher price for the good. Even if the price increase the demand will not decrease.

b) The amount consumer pay to the producer is calculated below:

At equilibrium the demand and the supply are equal Qd = Qs. 100 -P = -20 +P

= 120 = 2P

P = 120 / 2

= 60.

The price consumer pay to the producer is 60. The answer is "A".